If the average new home costs $145,000.00 today, how much will it cost in 20 years if the price increases by 3% each year?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=145,000*(1.03)^20

=145,000*1.80611123

=$261886.13(Approx).